# CityGaussianV2: Efficient and Geometrically Accurate Reconstruction for Large-Scale Scenes

### v1

@@ Appendix C More Implementation Details
For primitives and data partitioning, as well as parallel tuning, we follow the default parameter setting of CityGaussian (Liu et al., 2024) on both aerial view and street view of MatrixCity dataset. On GauU-Scene, we use SSIM threshold $\epsilon$ of 0.05 and default foreground range for contraction, i.e. the central 1/3 area of the scene. The Residence scene of GauU-Scene is divided into $4\times 2$ blocks, while Russian Building and Modern Building scenes are divided into $3\times 3$ blocks. When fine-tuning on GauU-Scene, the learning rate of position is reduced by 60%, while that of scaling is empirically reduced by 20%, as suggested in Liu et al. (2024). For vectree quantization, we set the codebook size to 8192 and the quantization ratio to 0.4.
@@ Appendix D Discussion
While our method successfully delivers favorable efficiency and accurate geometry reconstruction for large-scale scenes, we also want to discuss its limitations: First, for mesh extraction, we follow 2DGS to use a TSDF-based marching cube strategy. However, TSDF fusion struggles to balance thin structures’ quality and completeness, as figured out by Yu et al. (2024c). Developing a strategy that leverages the properties of Gaussian primitives alongside advanced mesh extraction algorithms could help address this issue. Secondly, although our compression strategy significantly enhances rendering speed, it still lags behind CityGS. Future work should explore deeper optimizations of rasterizers, such as those proposed by Feng et al. (2024), or the integration of Level of Detail (LoD) techniques Kerbl et al. (2024); Ren et al. (2024).

### v2

@@ Appendix C More Implementation Details
For primitives and data partitioning, as well as parallel tuning, we follow the default parameter setting of CityGaussian (Liu et al., 2024) on both aerial view and street view of MatrixCity dataset. To be specific, it applies lower learning rates during tuning compared to pertaining, and the street view is trained with a significantly lower learning rate and longer densification interval due to its extreme view sparsity (Zhou et al., 2024). On GauU-Scene, we use SSIM threshold $\epsilon$ of 0.05 and default foreground range for contraction, i.e. the central 1/3 area of the scene. The Residence scene of GauU-Scene is divided into $4\times 2$ blocks, while Russian Building and Modern Building scenes are divided into $3\times 3$ blocks. When fine-tuning on GauU-Scene, the learning rate of position is reduced by 60%, while that of scaling is empirically reduced by 20%, as suggested in Liu et al. (2024). For vectree quantization, we set the codebook size to 8192 and the quantization ratio to 0.4.
@@ Appendix D Discussion
While our method successfully delivers favorable efficiency and accurate geometry reconstruction for large-scale scenes, we also want to discuss its limitations: Firstly, this paper evaluates on the GauUScene and MatrixCity, which feature compensated or ideally constant lighting conditions. Nevertheless, we trust that the consideration of illumination variance and incorporating techniques like decoupled appearance modeling would be helpful for the model’s adaptability. Secondly, for mesh extraction, occlusion and lack of observation hinder reconstruction of some road surfaces and building facades. Additionally, TSDF fusion struggles with thin structures, such as spires shown in Fig. 11. Applying more efficient training strategies and advanced mesh extraction algorithms could address these issues. Thirdly, although our compression strategy significantly enhances rendering speed, it still lags behind CityGS even when sharing similar Gaussian counts. Extensive experiments in Tab. 4 validate this conclusion. Future work should explore deeper optimizations of rasterizers, such as those proposed by Feng et al. (2024), or the integration of Level of Detail (LoD) techniques Kerbl et al. (2024); Ren et al. (2024).